Roots of the equation 2 sin^2(A) + sin^2(2A) = 2 are - (a) p/6 (b) p/4 (c) p/3 (d) p/2 where p is for pi...

Grade:12th pass

2 Answers

jitender lakhanpal
10 years ago

hi

 

in these type of questions it is better to check from the options because there are only two terms 2 sin^2(A) + sin^2(2A)               put A = PIE / 2  you will see that this value satisfies the equation.

 

and for subjective take the second term sin^2(2A)  = 4 sin^2A * cos^2 A= 4 sin^2A (1-sin^2)...BY TRIGONOMETRIC IDENTITY

so the equation will become

4sin^2(A) - sin^4(A) + 2 sin^2 (A) = 2

put     sin^2A = t    you will get the quadratic equation 

then solve it for sin (A) you will get the answer

 

We are  here to help you in your IIT JEE preparation.

Now you can win by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar : Click here to download the toolbar.. 

 

Thanks and Regards

Jitender

(askIITians expert)

Himanshu Mohan
10 years ago

2sin^2(A)+sin^2(2A)=2

2sin^2(A)+4sin^2(A)cos^2(A)=2

2sin^2(A)+4sin^2(A){1-sin^2(A)}=2

sin^2(A)+2sin^2(A){1-sin^2(A)=1

2sin^4(A)-3sin^2(A)+1=0

sin^2(A)=3/4+_(9-8)^1/2/4

sin^2(A)=1/2 or1

sin(A)=+_(1/2)^1/2 Or +_1

A=P/6 OrP/2